# envname-P envname#1

### v2

@@ 1 Introduction
Large language models (LLMs) have emerged as a cornerstone of modern artificial intelligence, driving groundbreaking advancements in diverse domains such as machine translation, commonsense reasoning, and mathematical problem-solving. Despite their impressive capabilities, fine-tuning these models for downstream tasks introduces significant challenges [qiu2020pre, raffel2020exploring, rozière2024codellamaopenfoundation], primarily due to the substantial memory overhead required to store optimizer states, gradients, and intermediate activations. These memory footprints are particularly acute in resource-constrained environments, where the full-parameter fine-tuning of billion-scale models often exceeds the available hardware capacity, thereby impeding the practical deployment and scalability of LLMs [zhang2022opt, xu2023parameter, dubey2024llama].
@@
To address these challenges, parameter-efficient fine-tuning (PEFT) methods, such as Low-Rank Adaptation (LoRA) [hu2021lora] and its variants [liu2024dora, dettmers2023qloraefficientfinetuningquantized], have gained considerable attention. These methods freeze the majority of pre-trained parameters and optimize only small, low-rank matrices integrated into transformer layers, leading to substantial reductions in memory usage. However, while these approaches are highly memory-efficient, they inherently constrain the model’s adaptability. By optimizing only a sparse subset of parameters, such methods often lead to suboptimal performance compared to full-parameter fine-tuning, as task-specific features crucial to downstream performance may reside within the frozen portions of the network [lialin2023relorahighranktraininglowrank, zhang2024scaling].
@@
Recent research has explored block-coordinate descent (BCD) [tseng2001convergence, wright2015coordinate, xu2015block], a classical strategy for high-dimensional optimization, as a promising alternative for fine-tuning LLMs, resulting in the layer-wise optimization methods such as BAdam [luo2024badam], HIFT [liu2024hifthierarchicalparameterfinetuning], LIFT [zhu2024lift], LISA [pan2024lisa], and BlockLLM [ramesh2024blockllm]. Unlike LoRA, these layer-wise methods optimize transformer blocks sequentially while keeping the remaining layers frozen. By iteratively updating all transformer blocks, layer-wise optimization effectively enables full-parameter updates, thereby preserving the expressive capacity of the original model. Empirical studies demonstrate that layer-wise optimization consistently outperforms LoRA in performance [luo2024badam, pan2024lisa]. Furthermore, by skipping gradient computations for frozen layers, layer-wise optimization eliminates the need to store their intermediate activations, offering greater memory efficiency than LoRA.
@@
Motivating questions. All existing layer-wise LLM approaches [luo2024badam, liu2024hifthierarchicalparameterfinetuning, zhu2024lift, pan2024lisa, ramesh2024blockllm] adopt the traditional BCD optimization paradigm, in which the model weights are divided into manageable blocks and updated iteratively. Although BCD provides a structured framework for handling the high dimensionality of LLM weights, fully realizing its potential in practical LLM fine-tuning requires addressing the following fundamental open questions.
@@
The choice of partitioning strategy can significantly impact optimization efficiency and convergence. Current methods [luo2024badam, liu2024hifthierarchicalparameterfinetuning, zhu2024lift, pan2024lisa, ramesh2024blockllm] treat transformer layers as homogeneous units, overlooking the differing significance of internal modules—such as multi-head attention, feed-forward networks, and normalization layers—within each layer. By uniformly updating all parameters in a layer, these methods risk over-adapting less impactful modules while under-training critical ones, resulting in suboptimal optimization performance. Thus, more effective partitioning strategies must be explored.
@@
The performance of layer-wise methods is highly influenced by the block sampling strategies. However, current approaches primarily rely on cyclic [luo2024badam] or uniform [liu2024hifthierarchicalparameterfinetuning] sampling patterns, which overlook the varying importance across layers. LISA [pan2024lisa] keeps the Embedding and LLM head layers active but assigns equal sampling probabilities to all transformer layers, ignoring their varying significance. BlockLLM [ramesh2024blockllm] prioritize more frequent updates to critical layers but ultimately focus on a small set of fixed blocks, failing to adequately explore the remaining layers. An effective sampling strategy should maintain a balance between thorough exploration of the parameter space and efficient exploitation of the most promising optimization directions to ensure rapid convergence.
@@
Theoretical convergence of layer-wise LLM optimization remains understudied. Traditional BCD optimization literature typically focuses on block-wise GD/SGD convergence [wright2015coordinate, xu2013block, hong2017iteration]. Recent works [luo2024badam, zhou2020randomized, pan2024lisa] provide convergence guarantees under restrictive assumptions, such as noiseless gradients or single updates per sampled block. These analyses fail to reflect practical layer-wise LLM optimization, where Adam optimizer, stochastic gradients, and multiple updates per block are standard practice. A rigorous convergence analysis under these realistic conditions is essential for broader adoption of layer-wise optimization methods.
@@
    Module-wise optimization. We define a module as a matrix parameter within a transformer layer associated with weight gradients. Empirically, we observe that internal modules within transformer layers exhibit heterogeneous importance. Theoretically, we demonstrate that decomposing each layer into smaller modules preserves more information in gradient, thus motivating module-wise optimization. This addresses Question 1. Furthermore, we find this fine-grained update strategy eliminates the need to load a full layer into memory, making it more memory-efficient than layer-wise LLM optimization approaches [luo2024badam, liu2024hifthierarchicalparameterfinetuning, zhu2024lift, pan2024lisa, ramesh2024blockllm].
@@
Table 1: Comparison of fine-tuning methods on LLaMA3-8B across eight commonsense reasoning tasks. The ”ChatGPT” row presents results from ChatGPT obtained using Zero-shot CoT [wei2022chain] with the GPT-3.5-turbo API. "Hella." refers to the HellaSwag dataset [zellers2019hellaswag], and "Wino." refers to the Winogrande dataset [sakaguchi2021winogrande]. Memory usage is reported without the application of additional memory-saving techniques such as gradient checkpointing or flash attention. The symbol $\delta$ denotes the proportion of parameters updated in each training iteration.
@@
    Fine-tuning. We evaluated MISA on different LLMs across three benchmarks: Commonsense Reasoning [hu2023llm], Math Reasoning [hu2023llm], and Instruction Following, encompassing a total of 16 datasets. We compared MISA against PEFT and layer-wise optimization methods, including LoRA [hu2021lora], DoRA [liu2024dora], BAdam [luo2024badam], and LISA [pan2024lisa]. Under comparable memory constraints, MISA outperformed all baselines. An illustration of MISA’s superiority in fine-tuning tasks is provided in Table 1, with more detailed results presented in Section 5.
@@
    Pre-training. We trained the LLaMA2 130M and 350M variant [lialin2023relora] on the C4 dataset [raffel2020exploring]. In 350M model training, MISA achieved a perplexity of 22.11 after 2.7B training tokens, significantly outperforming GaLore’s 24.34 [zhao2024galore] , and approaching Adam’s 21.3.
@@ 2 Related Work
Parameter-efficient fine-tuning. Adapter-based methods introduce new modules (e.g., fully connected layers) into the backbone model and update only these modules during training [pfeiffer2020mad, houlsby2019parameter, he2021towards, wang2022adamix, he2022sparseadapter], incurring additional computational costs during inference. Prompt-based methods prepend random soft tokens to the input (typically as a prefix) and train their embeddings while keeping the backbone LLM frozen [li2021prefix, han2022ptr, zhong2021factual, qin2021learning], also introducing inference overhead.
@@
LoRA and its variants, including QLoRA [dettmers2024qlora], DoRA [liu2024dora], and AdaLoRA [zhang2023adaloraadaptivebudgetallocation], have become state-of-the-art PEFT methods. LoRA injects low-rank matrices into linear layers to estimate weight updates while keeping the backbone LLM parameters frozen. Unlike adapter- and prompt-based methods, LoRA has no additional inference overhead, as the low-rank matrices can be merged into the model weights. To address the limitations of the low-rank structure, recent methods such as ReLoRA [lialin2023relora], MoRA [jiang2024mora], and HiRA [anonymous2025hira] have explored high-rank updates based on LoRA.
@@
Memory-efficient optimization. LOMO [lv2023full] eliminates optimizer and gradient memory costs during training with standard SGD. AdaLOMO [lv2023adalomo] extends LOMO by integrating a fused backward operation with an adaptive learning rate per parameter . GaLore [zhao2024galore] enhances memory efficiency by projecting gradients into a low-rank subspace. Quantization techniques are widely applied to model parameters, gradients, optimizer states, and intermediate activations [dettmers20218, li2309memory, deepseekai2024deepseekv3technicalreport]. Adam-mini [zhang2024adamminiusefewerlearning] reduces memory usage by optimizing learning rate allocation in Adam.
@@
Layer-wise optimization. Inspired by classical Block Coordinate Descent (BCD), layer-wise optimization has emerged as a promising approach for fine-tuning large language models. Methods such as LIFT [zhu2024lift], HIFT [liu2024hifthierarchicalparameterfinetuning], and BAdam [luo2024badam] update LLM layers sequentially, achieving performance comparable to parameter-efficient fine-tuning (PEFT) methods while significantly reducing computational costs. LISA [pan2024lisa] identified a skewed weight-norm distribution across layers in LoRA and introduced a randomized layer update strategy, keeping the embedding and language modeling head layers activated. Similarly, BlockLLM [ramesh2024blockllm] prioritizes frequent updates to layers with larger gradient norms. However, none of these approaches incorporate module-wise importance sampling.
@@
Varying importance of modules in LLM fine-tuning. Previous studies have shown that different layers contribute unequally to LLM performance. LayerSkip [Elhoushi_2024], LayerDrop [fan2019reducingtransformerdepthdemand], and Layer-wise Model Pruning [fan2021layerwisemodelpruningbased] demonstrated that the overparameterization of pre-trained models leads to layer redundancy, enabling high performance to be maintained even after removing certain layers. IST [yao2024layerwiseimportancemattersmemory] and ILA [shi2024understandinglayersignificancellm] further highlighted that layer importance varies during LLM fine-tuning. Additionally, studies on PEFT have revealed that different modules within transformer blocks exhibit varying ranks [biderman2024lora, zhang2023adaloraadaptivebudgetallocation]. OwLore[li2024owloreoutlierweighedlayerwisesampled] identifies layer importance by detecting outliers in the model weights. However, these insights have not yet motivate module-wise importance sampling in LLM optimization.
@@ 3.3 Partition the Weights into Fine-Grained Modules Instead of Coarse Layers
Current approaches [luo2024badam, liu2024hifthierarchicalparameterfinetuning, zhu2024lift, pan2024lisa, ramesh2024blockllm] often treat transformer layers as homogeneous structures, neglecting the distinct roles and contributions of their internal components—such as multi-head attention mechanisms, feed-forward networks, and normalization layers. As shown in Fig. 1, different modules within the same transformer layer exhibit significantly varying gradient norms, underscoring their differing levels of importance. Uniformly updating all modules within a layer risks over-adapting less critical components while under-training more essential ones, ultimately leading to suboptimal performance.
@@ 3.4 Module-wise Importance Sampling (MISA) Method
LoRA[hu2021lora] IST[yao2024layerwiseimportancemattersmemory] GaLore[zhao2024galore] OwLore[li2024owloreoutlierweighedlayerwisesampled] BAdam[luo2024badam] LISA[pan2024lisa] MISA
@@
  • : GaLore’s convergence proof[zhao2024galore] holds only when the gradient exhibits a highly specialized structure, and BAdam’s analysis[luo2024badam] applies exclusively to the full‑gradient regime. Their settings in the theoretical guarantees rely on restrictive assumptions.
@@ 4 Convergence Analysis
Limitations in existing analysis. Conventional BCD methods focus on the convergence of block-wise GD or SGD [wright2015coordinate, xu2013block, hong2017iteration], but they do not generalize to the Adam optimizer. More recent studies [luo2024badam, zhou2020randomized, pan2024lisa] establish convergence results for layer-wise Adam, but only under restrictive conditions—such as assuming noiseless gradients or allowing just a single update per sampled block. These assumptions, however, diverge significantly from practical LLM training, where Adam is used with stochastic gradients and multiple updates are performed for each sampled block.
@@
According to Theorem 1, MISA converges for any number $T$ of inner local block updates, which constitutes a key advantage of our analysis. In contrast, the analyses in [wright2015coordinate, xu2013block, zhou2020randomized] are limited to the case $T=1$, while the analysis in [razaviyayn2013unified, chen2025memory] requires $T$ to be sufficiently large to ensure each block is adequately updated.
@@ 5.1 Commonsense Reasoning
Settings. To validate the efficiency of MISA, we fine-tuned the LLaMA3-8B [dubey2024llama] and Qwen2.5-7B [qwen2] models on the Commonsense Reasoning tasks. The dataset consists of eight subsets: BoolQ [clark2019boolq], PIQA [bisk2020piqa], SIQA [sap2019socialiqa], HellaSwag [zellers2019hellaswag], WinoGrande [sakaguchi2021winogrande], ARC [clark2018think], and OBQA [mihaylov2018can]. Following the settings in [hu2023llm], we combined the training data from all eight tasks into a single training set for fine-tuning and then evaluated each model separately on the eight test sets. The baselines included LoRA [hu2021lora], DoRA [liu2024dora], and BCD-based methods, namely LISA [pan2024lisa] and BAdam [luo2024badam]. We also reported GPU memory consumption for each method to ensure a thorough comparison. No additional memory-saving techniques, such as checkpointing [chen2016trainingdeepnetssublinear] or flash attention [dao2023flashattention2], were employed.
Table 3: This table compares fine-tuning methods on Qwen2.5-7B models across eight commonsense reasoning tasks. "Hella." refers to HellaSwag [zellers2019hellaswag], and "Wino." to Winogrande [sakaguchi2021winogrande]. The first row shows ChatGPT results from Zero-shot CoT [wei2022chain] using the GPT-3.5-turbo API. Refer to Table 1 for results of LLaMA3-8B.

### v3

@@ 1 Introduction
Large language models (LLMs) have emerged as a cornerstone of modern artificial intelligence, driving groundbreaking advancements in diverse domains such as machine translation, commonsense reasoning, and mathematical problem-solving. Despite their impressive capabilities, fine-tuning these models for downstream tasks introduces significant challenges [45, 46, rozière2024codellamaopenfoundation], primarily due to the substantial memory overhead required to store optimizer states, gradients, and intermediate activations. These memory footprints are particularly acute in resource-constrained environments, where the full-parameter fine-tuning of billion-scale models often exceeds the available hardware capacity, thereby impeding the practical deployment and scalability of LLMs [68, 59, 14].
@@
To address these challenges, parameter-efficient fine-tuning (PEFT) methods, such as Low-Rank Adaptation (LoRA) [24] and its variants [35, 12], have gained considerable attention. These methods freeze the majority of pre-trained parameters and optimize only small, low-rank matrices integrated into transformer layers, leading to substantial reductions in memory usage. However, while these approaches are highly memory-efficient, they inherently constrain the model’s adaptability. By optimizing only a sparse subset of parameters, such methods often lead to suboptimal performance compared to full-parameter fine-tuning, as task-specific features crucial to downstream performance may reside within the frozen portions of the network [33, 65].
@@
Recent research has explored block-coordinate descent (BCD) [54, 58, 61], a classical strategy for high-dimensional optimization, as a promising alternative for fine-tuning LLMs, resulting in the layer-wise optimization methods such as BAdam [37], HIFT [36], LIFT [74], LISA [41], and BlockLLM [47]. Unlike LoRA, these layer-wise methods optimize transformer blocks sequentially while keeping the remaining layers frozen. By iteratively updating all transformer blocks, layer-wise optimization effectively enables full-parameter updates, thereby preserving the expressive capacity of the original model. Empirical studies demonstrate that layer-wise optimization consistently outperforms LoRA in performance [37, 41]. Furthermore, by skipping gradient computations for frozen layers, layer-wise optimization eliminates the need to store their intermediate activations, offering greater memory efficiency than LoRA.
@@
Motivating questions. All existing layer-wise LLM approaches [37, 36, 74, 41, 47] adopt the traditional BCD optimization paradigm, in which the model weights are divided into manageable blocks and updated iteratively. Although BCD provides a structured framework for handling the high dimensionality of LLM weights, fully realizing its potential in practical LLM fine-tuning requires addressing the following fundamental open questions.
@@
The choice of partitioning strategy can significantly impact optimization efficiency and convergence. Current methods [37, 36, 74, 41, 47] treat transformer layers as homogeneous units, overlooking the differing significance of internal modules—such as multi-head attention, feed-forward networks, and normalization layers—within each layer. By uniformly updating all parameters in a layer, these methods risk over-adapting less impactful modules while under-training critical ones, resulting in suboptimal optimization performance. Thus, more effective partitioning strategies must be explored.
@@
The performance of layer-wise methods is highly influenced by the block sampling strategies. However, current approaches primarily rely on cyclic [37] or uniform [36] sampling patterns, which overlook the varying importance across layers. LISA [41] keeps the Embedding and LLM head layers active but assigns equal sampling probabilities to all transformer layers, ignoring their varying significance. BlockLLM [47] prioritize more frequent updates to critical layers but ultimately focus on a small set of fixed blocks, failing to adequately explore the remaining layers. An effective sampling strategy should maintain a balance between thorough exploration of the parameter space and efficient exploitation of the most promising optimization directions to ensure rapid convergence.
@@
Theoretical convergence of layer-wise LLM optimization remains understudied. Traditional BCD optimization literature typically focuses on block-wise GD/SGD convergence [58, 60, 22]. Recent works [37, 73, 41] provide convergence guarantees under restrictive assumptions, such as noiseless gradients or single updates per sampled block. These analyses fail to reflect practical layer-wise LLM optimization, where Adam optimizer, stochastic gradients, and multiple updates per block are standard practice. A rigorous convergence analysis under these realistic conditions is essential for broader adoption of layer-wise optimization methods.
@@
    Module-wise optimization. We define a module as a matrix parameter within a transformer layer associated with weight gradients. Empirically, we observe that internal modules within transformer layers exhibit heterogeneous importance. Theoretically, we demonstrate that decomposing each layer into smaller modules preserves more information in gradient, thus motivating module-wise optimization. This addresses Question 1. Furthermore, we find this fine-grained update strategy eliminates the need to load a full layer into memory, making it more memory-efficient than layer-wise LLM optimization approaches [37, 36, 74, 41, 47].
@@
Table 1: Comparison of fine-tuning methods on LLaMA3-8B across eight commonsense reasoning tasks. The ”ChatGPT” row presents results from ChatGPT obtained using Zero-shot CoT [57] with the GPT-3.5-turbo API. "Hella." refers to the HellaSwag dataset [64], and "Wino." refers to the Winogrande dataset [49]. Memory usage is reported without the application of additional memory-saving techniques such as gradient checkpointing or flash attention. The symbol $\delta$ denotes the proportion of parameters updated in each training iteration.
@@
    Fine-tuning. We evaluated MISA on different LLMs across three benchmarks: Commonsense Reasoning [25], Math Reasoning [25], and Instruction Following, encompassing a total of 16 datasets. We compared MISA against PEFT and layer-wise optimization methods, including LoRA [24], DoRA [35], BAdam [37], and LISA [41]. Under comparable memory constraints, MISA outperformed all baselines. An illustration of MISA’s superiority in fine-tuning tasks is provided in Table 1, with more detailed results presented in Section 5.
@@
    Pre-training. We trained the LLaMA2 130M and 350M variant [32] on the C4 dataset [46]. In 350M model training, MISA achieved a perplexity of 22.11 after 2.7B training tokens, significantly outperforming GaLore’s 24.34 [70] , and approaching Adam’s 21.3.
@@ 2 Related Work
Parameter-efficient fine-tuning. Adapter-based methods introduce new modules (e.g., fully connected layers) into the backbone model and update only these modules during training [43, 23, 19, 55, 20], incurring additional computational costs during inference. Prompt-based methods prepend random soft tokens to the input (typically as a prefix) and train their embeddings while keeping the backbone LLM frozen [31, 18, 72, 44], also introducing inference overhead.
@@
LoRA and its variants, including QLoRA [13], DoRA [35], and AdaLoRA [67], have become state-of-the-art PEFT methods. LoRA injects low-rank matrices into linear layers to estimate weight updates while keeping the backbone LLM parameters frozen. Unlike adapter- and prompt-based methods, LoRA has no additional inference overhead, as the low-rank matrices can be merged into the model weights. To address the limitations of the low-rank structure, recent methods such as ReLoRA [32], MoRA [27], and HiRA [1] have explored high-rank updates based on LoRA.
@@
Memory-efficient optimization. LOMO [39] eliminates optimizer and gradient memory costs during training with standard SGD. AdaLOMO [38] extends LOMO by integrating a fused backward operation with an adaptive learning rate per parameter . GaLore [70] enhances memory efficiency by projecting gradients into a low-rank subspace. Quantization techniques are widely applied to model parameters, gradients, optimizer states, and intermediate activations [11, 29, 10]. Adam-mini [69] reduces memory usage by optimizing learning rate allocation in Adam.
@@
Layer-wise optimization. Inspired by classical Block Coordinate Descent (BCD), layer-wise optimization has emerged as a promising approach for fine-tuning large language models. Methods such as LIFT [74], HIFT [36], and BAdam [37] update LLM layers sequentially, achieving performance comparable to parameter-efficient fine-tuning (PEFT) methods while significantly reducing computational costs. LISA [41] identified a skewed weight-norm distribution across layers in LoRA and introduced a randomized layer update strategy, keeping the embedding and language modeling head layers activated. Similarly, BlockLLM [47] prioritizes frequent updates to layers with larger gradient norms. However, none of these approaches incorporate module-wise importance sampling.
@@
Varying importance of modules in LLM fine-tuning. Previous studies have shown that different layers contribute unequally to LLM performance. LayerSkip [15], LayerDrop [16], and Layer-wise Model Pruning [17] demonstrated that the overparameterization of pre-trained models leads to layer redundancy, enabling high performance to be maintained even after removing certain layers. IST [63] and ILA [51] further highlighted that layer importance varies during LLM fine-tuning. Additionally, studies on PEFT have revealed that different modules within transformer blocks exhibit varying ranks [2, 67]. OwLore[30] identifies layer importance by detecting outliers in the model weights. However, these insights have not yet motivate module-wise importance sampling in LLM optimization.
@@ 3.3 Partition the Weights into Fine-Grained Modules Instead of Coarse Layers
Current approaches [37, 36, 74, 41, 47] often treat transformer layers as homogeneous structures, neglecting the distinct roles and contributions of their internal components—such as multi-head attention mechanisms, feed-forward networks, and normalization layers. As shown in Fig. 1, different modules within the same transformer layer exhibit significantly varying gradient norms, underscoring their differing levels of importance. Uniformly updating all modules within a layer risks over-adapting less critical components while under-training more essential ones, ultimately leading to suboptimal performance.
@@ 3.4 Module-wise Importance Sampling (MISA) Method
LoRA[24] IST[63] GaLore[70] OwLore[30] BAdam[37] LISA[41] MISA
@@
  • : GaLore’s convergence proof[70] holds only when the gradient exhibits a highly specialized structure, and BAdam’s analysis[37] applies exclusively to the full‑gradient regime. Their settings in the theoretical guarantees rely on restrictive assumptions.
@@ 4 Convergence Analysis
Limitations in existing analysis. Conventional BCD methods focus on the convergence of block-wise GD or SGD [58, 60, 22], but they do not generalize to the Adam optimizer. More recent studies [37, 73, 41] establish convergence results for layer-wise Adam, but only under restrictive conditions—such as assuming noiseless gradients or allowing just a single update per sampled block. These assumptions, however, diverge significantly from practical LLM training, where Adam is used with stochastic gradients and multiple updates are performed for each sampled block.
@@
According to Theorem 1, MISA converges for any number $T$ of inner local block updates, which constitutes a key advantage of our analysis. In contrast, the analyses in [58, 60, 73] are limited to the case $T=1$, while the analysis in [48, 5] requires $T$ to be sufficiently large to ensure each block is adequately updated.
@@ 5.1 Commonsense Reasoning
Settings. To validate the efficiency of MISA, we fine-tuned the LLaMA3-8B [14] and Qwen2.5-7B [62] models on the Commonsense Reasoning tasks. The dataset consists of eight subsets: BoolQ [6], PIQA [3], SIQA [50], HellaSwag [64], WinoGrande [49], ARC [7], and OBQA [40]. Following the settings in [25], we combined the training data from all eight tasks into a single training set for fine-tuning and then evaluated each model separately on the eight test sets. The baselines included LoRA [24], DoRA [35], and BCD-based methods, namely LISA [41] and BAdam [37]. We also reported GPU memory consumption for each method to ensure a thorough comparison. No additional memory-saving techniques, such as checkpointing [4] or flash attention [9], were employed.
Table 3: This table compares fine-tuning methods on Qwen2.5-7B models across eight commonsense reasoning tasks. "Hella." refers to HellaSwag [64], and "Wino." to Winogrande [49]. The first row shows ChatGPT results from Zero-shot CoT [57] using the GPT-3.5-turbo API. Refer to Table 1 for results of LLaMA3-8B.